# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ V-D Problem Formulation
$\mathrm{s.t.}\ \$$\pi_{\mathbbm{j}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}},\quad\forall\,u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}}$ (22a)
$x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}=0,\text{if~}s^{\mathrm{On}}_{\mathbbm{i}}\notin\bm{C}_{\mathbbm{i}},~\forall\,u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}}$ (22b)
$\sum_{u^{\mathrm{On}}_{\mathbbm{j}}\in{\mathcal{U}^{\mathrm{On}}}}x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}\leq G_{\mathbbm{i}},\quad\forall\,s^{\mathrm{On}}_{\mathbbm{i}}\in\mathcal{S}^{\mathrm{On}}$ (22c)
where $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$ represents the joint scheduling strategy; constraint (22a) enforces that each SD chooses either local execution or exactly one SP; constraint (22b) encodes the service‐coverage feasibility condition: any SD located outside the service region of SP $u_{j}$, is ineligible for assignment. constraint (22c) ensures that the number of SDs served by each SP does not exceed its maximum access $G_{\mathbbm{i}}$.
@@
1Input: $\mathcal{U}^{\mathrm{On}}$, $\mathcal{S}^{\mathrm{On}}$, and max iterations $T_{\max}$.
2[2pt]
3Initialization: Randomly construct an initial feasible profile $\bm{\pi}^{(0)}=(\pi^{(0)}_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$ such that constraints (22a)–(22c) are satisfied, set $\mathbbm{t}\leftarrow 0$
4while not converged and $\mathbbm{t}<T_{\max}$ do
5
6 For each SP $s^{\mathrm{On}}_{\mathbbm{i}}\in\mathcal{S}^{\mathrm{On}}$, compute $y_{\mathbbm{i}}(\bm{\pi}^{(\mathbbm{t})}),\mathbbm{x}_{\mathbbm{i}}(\bm{\pi}^{(\mathbbm{t})})$, and $n_{\mathbbm{i}}(\bm{\pi}^{(\mathbbm{t})})$.
7 generate a random permutation of $\mathcal{U}^{\mathrm{On}}$.
8 for each SD $u^{\mathrm{On}}_{\mathbbm{j}}$ in the permutation do
9     let $a^{\mathrm{cur}}_{\mathbbm{j}}=\pi^{(\mathbbm{t})}_{\mathbbm{j}}$; for all $\mathbbm{i}$, obtain $y^{-{\mathbbm{j}}}_{\mathbbm{i}}$, $\mathbbm{x}^{-{\mathbbm{j}}}_{\mathbbm{i}}$, $n^{-{\mathbbm{j}}}_{\mathbbm{i}}$
10    build $\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}$ according to (23).
11    compute current utility $U_{\mathbbm{j}}(a^{\mathrm{cur}}_{\mathbbm{j}}\mid\bm{\pi}^{(\mathbbm{t})}_{-\mathbbm{j}})$ via (19);
12     for each $a\in\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}$, form hypothetical loads $y^{(\mathbbm{j}\rightarrow a)}_{\mathbbm{i}}$, $\mathbbm{x}^{(\mathbbm{j}\rightarrow a)}_{\mathbbm{i}}$ and evaluate $U_{\mathbbm{j}}(a\mid\bm{\pi}^{(\mathbbm{t})}_{-\mathbbm{j}})$;
13    let $a^{\mathrm{best}}_{\mathbbm{j}}\in\arg\max_{a\in\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}}U_{\mathbbm{j}}(a\mid\bm{\pi}^{(\mathbbm{t})}_{-\mathbbm{j}})$.
14    if $U_{\mathbbm{j}}(a^{\mathrm{best}}_{\mathbbm{j}}\mid\bm{\pi}^{(\mathbbm{t})}_{-\mathbbm{j}})-U_{\mathbbm{j}}(a^{\mathrm{cur}}_{\mathbbm{j}}\mid\bm{\pi}^{(\mathbbm{t})}_{-\mathbbm{j}})>\varepsilon$ then
15        set $\pi^{(\mathbbm{t}+1)}_{\mathbbm{j}}\leftarrow a^{\mathrm{best}}_{\mathbbm{j}}$
16    else
17        set $\pi^{(\mathbbm{t}+1)}_{\mathbbm{j}}\leftarrow\pi^{(\mathbbm{t})}_{\mathbbm{j}}$
18
19
20 if no SD updates its action in this pass, declare convergence;
21  $\mathbbm{t}\leftarrow\mathbbm{t}+1$.
22
Output: NE profile $\bm{\pi}^{\star}$ (final $\bm{\pi}^{(\mathbbm{t})}$) and the associated allocation.
Algorithm 3 Proposed PG-BRD
V-E Potential-Guided Best-Response Dynamics (PG-BRD)
We consider an online ordinal-potential game among SDs and SPs under the hard coverage and subcarrier-capacity constraints in (22b)–(22c). Each SD myopically updates its decision to maximize its own utility, while the underlying ordinal potential $\Phi(\bm{\pi})$ guarantees convergence to a pure-strategy NE.
@@ V-E Potential-Guided Best-Response Dynamics (PG-BRD)
Step 1. Random-feasible initialization. We first construct a random feasible association profile $\bm{\pi}^{(0)}\in\{0,1,\ldots,|\mathcal{S}^{\mathrm{On}}|\}^{|\mathcal{U}^{\mathrm{On}}|}$, where $\pi_{\mathbbm{j}}=0$ corresponds to local execution and $\pi_{\mathbbm{j}}=\mathbbm{i}>0$ corresponds to transferring task data to SP $s^{\mathrm{On}}_{\mathbbm{i}}$. For each SD $u^{\mathrm{On}}_{\mathbbm{j}}$, we randomly select an action from $\{0\}\cup\bm{C}_{\mathbbm{j}}$ such that the resulting profile satisfies the coverage/access constraint (22b) and the subcarrier constraint (22c). This random-feasible initialization avoids starting from an infeasible point and provides a diverse baseline for the subsequent best-response updates.
@@
Step 2. Feasible-action enumeration under coverage and subcarrier caps. Given the current profile $\bm{\pi}$, we compute for each SP $s^{\mathrm{On}}_{\mathbbm{i}}$ the effective congestion load $y_{\mathbbm{i}}(\bm{\pi})$ and the physical computing load $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$ as in (20), as well as the number of attached SDs $n_{\mathbbm{i}}(\bm{\pi})$. When updating a particular SD $u^{\mathrm{On}}_{\mathbbm{j}}$, we first remove its current contribution, obtaining $y^{-{\mathbbm{j}}}_{\mathbbm{i}}$, $\mathbbm{x}^{-{\mathbbm{j}}}_{\mathbbm{i}}$, and $n^{-{\mathbbm{j}}}_{\mathbbm{i}}$. Based on these “others-fixed” loads, we construct the SD’s feasible action set $\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}$ as in (23): local execution $a=0$ is always feasible, while an offloading action $a=\mathbbm{i}>0$ is feasible only if $s^{\mathrm{On}}_{\mathbbm{i}}\in\bm{C}_{\mathbbm{j}}$ and the resulting number of served SDs does not exceed the subcarrier cap, i.e., $n^{-{\mathbbm{j}}}_{\mathbbm{i}}+1\leq G_{\mathbbm{i}}$. This step hard-prunes any action that would violate coverage or SP capacity constraints.
$\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}=\Big\{a\in\{0\}\cup\mathcal{S}^{\mathrm{On}}\ \Big|\ a=0\ \text{or}\ \big(a=\mathbbm{i}>0,\ s^{\mathrm{On}}_{\mathbbm{i}}\in\bm{C}_{\mathbbm{j}},\ n^{-{\mathbbm{j}}}_{\mathbbm{i}}+1\leq G_{\mathbbm{i}}\big)\Big\},$ (23)
@@
Step 3. Asynchronous best response under the ordinal potential. We adopt asynchronous updates: at each iteration, SDs are visited in a random order. For a given SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and current profile $\bm{\pi}$, we first evaluate its current utility $U_{\mathbbm{j}}(\pi_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})$ according to the unified form in (19). Then, for each feasible action $a\in\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}$, we form the hypothetical congestion and load variables and compute the corresponding utility $U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$. The SD selects a best response $a^{\mathrm{best}}_{\mathbbm{j}}\in\arg\max_{a\in\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}}U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$. To avoid oscillations caused by marginal improvements, we employ an acceptance threshold $\varepsilon>0$: the SD updates its decision only if
$U_{\mathbbm{j}}(a^{\mathrm{best}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})-U_{\mathbbm{j}}(\pi_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})>\varepsilon.$
Otherwise, it keeps $\pi_{\mathbbm{j}}$ unchanged. Because the game admits the ordinal potential $\Phi(\bm{\pi})$, each accepted update strictly increases $\Phi(\bm{\pi})$.
@@
Step 4. Convergence and NE detection. We repeatedly perform feasible-action enumeration and asynchronous best responses by cycling over all SDs. When a full pass over $\mathcal{U}^{\mathrm{On}}$ yields no accepted update, no SD can further improve its utility by more than $\varepsilon$ via a unilateral deviation, and the algorithm has reached a pure-strategy NE $\bm{\pi}^{\star}$. The resulting equilibrium profile $\bm{\pi}^{\star}$ serves as a locally optimal solution to problem (22), balancing user-side benefits and SP-side costs under congestion-aware service scheduling.
@@ VI Evaluations
This section conducts extensive experiments to evaluate our FUSION, via testing both numerical and real-world datasets on various evaluation metrics..
@@
To ensure statistical reliability, all figures are obtained via Monte Carlo method, where each data point represents the average result over 100 independent trials.
@@ VI-A Benchmark Methods and Evaluation Metrics
We evaluate our FUSION from two complementary angles: (i) prediction performance (regarding offline stage) on service demand forecasting module (Sec. VI-B); and (ii) system-level performance of FUSION (Sec. VI-C).
VI-A1 Benchmarks and Metrics for Demand Forecasting (offline stage)
We first compare our Pro-LNN module against two widely used prediction models.
$\bullet$ LSTM-driven demand prediction[27]: A standard long short-term memory network applied to the same input sequences $\{\mathbf{z}_{i}\}$, trained to predict future ES demands. This baseline represents a strong discrete-time RNN.
$\bullet$ Transformer-driven demand prediction[28]: A temporal Transformer-based predictor with self-attention over the historical context window. This serves as a powerful high-capacity baseline, but typically requires more parameters and computing resources.
To achieve better evaluations, we consider the following criteria:
@@ VI-A1 Benchmarks and Metrics for Demand Forecasting (offline stage)
$\bullet$ Model footprint: We measure the total number of trainable parameters and the corresponding FP32 model size (in megabytes) of each predictor, which together characterize its memory and storage requirements on edge devices.
@@
$\bullet$ Robustness: To assess robustness under realistic noisy conditions, we compare the mean squared error (MSE) on clean test inputs with that on corrupted inputs containing randomly missing observations, additive perturbations, and occasional demand bursts, and report the resulting performance drop $\Delta\text{MSE}$.
VI-A2 System-Level Benchmarks and Metrics (For Comparing With Our FUSION)
We then compare FUSION with several benchmark methods that modify or simplify either stages, to highlight the benefits of prediction-driven contracting, time-window-aware routing, and potential-based online coordination.
@@ VI-A2 System-Level Benchmarks and Metrics (For Comparing With Our FUSION)
$\bullet$ Pure online scheduling (PurOnline)[13]: This baseline removes the offline stage completely, e.g., ESs and AgPs do not sign any offline contracts, and UAV locations are not pre-planned. All HuUs and MaUs directly enter the online market, where their tasks are scheduled only via the PG-BRD-based potential game over the current ES/AgP topology.
@@
$\bullet$ Two-stage without PG-based online scheduling (FUSION_NoPG): This variant keeps the same offline stage as FUSION, i.e., Off-AIC2 with eACO-VRP is still used to form offline AgP–ES contracts and pre-planned routes. However, the online stage abandons PG-BRD: HuU/MaU tasks are randomly assigned to feasible ESs/AgPs within coverage, subject to capacity and deadline constraints. This aims to isolate the contribution of the online potential game.
@@
$\bullet$ Two-stage with random offline contracts (FUSION_Random): Here, an online stage is identical to FUSION and still relies on PG-BRD for congestion-aware online scheduling. In contrast, offline stage no longer uses Off-AIC2 or eACO-VRP; instead, each AgP is randomly associated with a subset of ESs and moves along a randomly generated feasible route until its capacity is exhausted. This quantifies the gain brought by a well-designed economic and path-aware offline contracting.
@@
$\bullet$ Two-stage without ES spatio–temporal features (FUSION_NoST): This preserves the two-stage structure, but the offline auction ignores ES-specific time windows and travel-time constraints. AgP–ES pairs are determined only by simple price/volume trading without using eACO-VRP, while online stage still runs PG-BRD for task scheduling. This aims highlighting the benefits of explicitly modeling ES spatio–temporal characteristics during offline stage.
To provide a comprehensive performance assessment, we adopt the following metrics.
$\bullet$ Social welfare (SW): The sum of utilities of all ESs, the AgPs, and users across both stages. In the offline stage, SW accounts for ES profits, AgP utilities, and auctioneer surplus as in (1)–(3). Then, during offline stage, it is aligned with the potential value $\Phi(\bm{\pi})$, which combines SD-side benefits and SP-side costs under congestion.
$\bullet$ Delay incurred by interactions between SPs and SDs (DoI): This captures the total decision–making latency (in milliseconds) induced by control-message exchanges between ESs/AgPs (SPs) and HuUs/MaUs (SDs) across both stages. Following [29, 30, 31], we assume that the per–interaction delay lies in the range of $[1,15]$ ms, and obtain DoI by multiplying the total number of interactions with the corresponding per–interaction delay.
[…]
$\bullet$ Truthfulness and individual rationality of Off-AIC2: To empirically verify the theoretical guarantees of Off-AIC2, we conduct experiments in which ESs and AgPs are allowed to deviate from their truthful bids/asks. We measure the proportion of agents whose utilities increase under such deviations. An auction that is truthful and individually rational should ensure that (i) every winning ES attains a non-negative utility and cannot benefit from misreporting its bid, and (ii) every winning AgP receives a payment no lower than its declared ask and gains no advantage by inflating or deflating it.
@@
TABLE II: Robustness–complexity comparison of different forecasters on synthetic non-stationary series. The “corrupt” case injects 30% random missing values, Gaussian noise with $\sigma=0.05$, and random burst shocks.
@@
TABLE III: Prediction accuracy and model footprint on the UCI Electricity Load Diagrams 2011–2014 dataset (feeders MT_005 and MT_145, input length $T_{\mathrm{his}}=48$, prediction horizon $H=10$).

### v2

@@ V-C Problem Formulation
$\mathrm{s.t.}\ \$$\pi_{\mathbbm{j}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}},\quad\forall\,u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}},$ (22a)
$x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}=0,\ \ \text{if }s^{\mathrm{On}}_{\mathbbm{i}}\notin\bm{C}_{\mathbbm{j}},\quad\forall\,u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}},$ (22b)
$\sum_{u^{\mathrm{On}}_{\mathbbm{j}}\in{\mathcal{U}^{\mathrm{On}}}}x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}\leq G_{\mathbbm{i}},\quad\forall\,s^{\mathrm{On}}_{\mathbbm{i}}\in\mathcal{S}^{\mathrm{On}},$ (22c)
where $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$ denotes the joint assignment profile, with $\pi_{\mathbbm{j}}=0$ representing local task execution at $u^{\mathrm{On}}_{\mathbbm{j}}$ and $\pi_{\mathbbm{j}}=\mathbbm{i}$ indicating task offloading to SP $s^{\mathrm{On}}_{\mathbbm{i}}$. The binary variables $x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}$ are induced by $\bm{\pi}$ as defined in Sec. V-B (i.e., $x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}=1$ if $\pi_{\mathbbm{j}}=\mathbbm{i}$).
Constraint (22a) enforces each SD to choose only one execution option for its current task: either local processing ($\pi_{\mathbbm{j}}=0$) or offloading to a single SP ($\pi_{\mathbbm{j}}\in\mathcal{S}^{\mathrm{On}}$). Constraint (22b) captures the service-coverage feasibility: $\bm{C}_{\mathbbm{j}}$ denotes the set of SPs that can physically serve $u^{\mathrm{On}}_{\mathbbm{j}}$ (e.g., within communication range), and thus the association variable for an SD outside the coverage of $s^{\mathrm{On}}_{\mathbbm{i}}$ is forced to zero. Constraint (22c) limits the admission load at each SP, where $G_{\mathbbm{i}}$ is the maximum number of SDs that SP $s^{\mathrm{On}}_{\mathbbm{i}}$ can concurrently serve.
@@
V-D Potential-Guided Best-Response Dynamics (PG-BRD)
For brevity, we provide the complete pseudo-code of PG-BRD in Appx. D and summarize its four main steps below:
@@ V-D Potential-Guided Best-Response Dynamics (PG-BRD)
Step 1. Random-feasible initialization: We first construct a feasible initial association profile $\bm{\pi}^{(0)}\in\{0,1,\ldots,|\mathcal{S}^{\mathrm{On}}|\}^{|\mathcal{U}^{\mathrm{On}}|}$, where $\pi_{\mathbbm{j}}=0$ corresponds to local task execution at $u^{\mathrm{On}}_{\mathbbm{j}}$ and $\pi_{\mathbbm{j}}=\mathbbm{i}>0$ means that $u^{\mathrm{On}}_{\mathbbm{j}}$ offloads its task to SP $s^{\mathrm{On}}_{\mathbbm{i}}$. For each SD $u^{\mathrm{On}}_{\mathbbm{j}}$, we randomly select an action from the set $\{0\}\cup\bm{C}_{\mathbbm{j}}$ and only accept it if the resulting profile satisfies the coverage constraint (22b) and the subcarrier-capacity constraint (22c). This procedure ensures that the algorithm starts from a valid resource scheduling configuration with no constraint violations.
@@
Step 2. Feasible-action enumeration under coverage and subcarrier-capacity constraints: Given a current profile $\bm{\pi}$, we compute for each SP $s^{\mathrm{On}}_{\mathbbm{i}}$ the effective congestion load $y_{\mathbbm{i}}(\bm{\pi})$ and the physical computing load $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$ as in (20), as well as the number of currently served SDs $n_{\mathbbm{i}}(\bm{\pi})=\big|\{\,\mathbbm{j}\mid\pi_{\mathbbm{j}}=\mathbbm{i}\,\}\big|$. When it is SD $u^{\mathrm{On}}_{\mathbbm{j}}$’s turn to update its decision, we temporarily remove its own contribution and obtain the loads generated by all other SDs, denoted by $y^{-{\mathbbm{j}}}_{\mathbbm{i}}$, $\mathbbm{x}^{-{\mathbbm{j}}}_{\mathbbm{i}}$, and $n^{-{\mathbbm{j}}}_{\mathbbm{i}}$. These quantities represent the congestion level and computing load at each SP if $u^{\mathrm{On}}_{\mathbbm{j}}$ were not associated with any SP. Given these congestion and load levels generated by all other SDs, we construct the feasible action set for $u^{\mathrm{On}}_{\mathbbm{j}}$ as follows:
$\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}{=}\Big\{a\in\{0\}\cup\mathcal{S}^{\mathrm{On}}\hskip 0.0pt\Big|\hskip 0.0pta=0\ \text{or}\ \big(a=\mathbbm{i}>0,\ s^{\mathrm{On}}_{\mathbbm{i}}\in\bm{C}_{\mathbbm{j}},\ n^{-{\mathbbm{j}}}_{\mathbbm{i}}+1\leq G_{\mathbbm{i}}\big)\Big\},$
where $a=0$ corresponds to local task execution, and an offloading action $a=\mathbbm{i}>0$ is feasible only if SP $s^{\mathrm{On}}_{\mathbbm{i}}$ covers $u^{\mathrm{On}}_{\mathbbm{j}}$ and still has enough residual subcarrier-capacity to admit one more SD without violating its capacity $G_{\mathbbm{i}}$. This step explicitly removes any candidate action that would break either the coverage or the subcarrier-capacity constraints.
@@
Step 3. Asynchronous best response under the ordinal potential: In each iteration, SDs are visited one-by-one in a random order, and at most one SD is allowed to change its decision at a time. For a given SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and current profile $\bm{\pi}$, we first evaluate its current utility $U_{\mathbbm{j}}(\pi_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})$ using the unified expression in (19). Then, for each feasible action $a\in\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}$, we temporarily set $\pi_{\mathbbm{j}}=a$, recompute the congestion and delay quantities for this SD, and obtain the resulting utility $U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$. The SD then chooses a best response $a^{\mathrm{best}}_{\mathbbm{j}}\in\arg\max_{a\in\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}}U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$. To avoid frequent switches caused by marginal gains, we introduce an acceptance threshold $\varepsilon>0$ and let $u^{\mathrm{On}}_{\mathbbm{j}}$ update its decision only if $U_{\mathbbm{j}}(a^{\mathrm{best}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})-U_{\mathbbm{j}}(\pi_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})>\varepsilon$. Otherwise, it keeps its current action $\pi_{\mathbbm{j}}$ unchanged. Because the game admits the ordinal potential $\Phi(\bm{\pi})$, every accepted update in this step strictly increases $\Phi(\bm{\pi})$.
@@
Step 4. Convergence and resulting equilibrium: The above two steps (feasible-action enumeration and asynchronous best responses) are repeatedly applied by going over all SDs. When a complete pass over $\mathcal{U}^{\mathrm{On}}$ produces no accepted update, no SD can improve its utility by more than $\varepsilon$ through a unilateral change of its action, and thus a pure-strategy NE $\bm{\pi}^{\star}$ is reached.
@@ VI-A Benchmark Methods and Evaluation Metrics
We evaluate FUSION from two complementary angles: (i) prediction performance on service demand forecasting module (Sec. VI-B); and (ii) system-level performance (Sec. VI-C).
VI-A1 Benchmarks and Metrics for Demand Forecasting
We compare Pro-LNN against two widely used prediction models.
$\bullet$ LSTM-driven demand prediction[28]: A standard long short-term memory (LSTM) network is applied to the input sequences $\{\mathbf{z}_{i}\}$ and trained to predict future ES demands. This baseline represents a strong discrete-time RNN model.
$\bullet$ Transformer-driven demand prediction[29]: A temporal Transformer-based predictor with self-attention is used over the historical context window. This serves as a powerful baseline, but typically requires more computing resources.
Performance evaluations consider the following metrics:
@@ VI-A1 Benchmarks and Metrics for Demand Forecasting
$\bullet$ Model footprint: We measure the total number of trainable parameters and the corresponding 32-bit floating-point format (FP32, measured in megabytes) for each predictor, which together characterize its computing (memory/storage) overhead.
@@
$\bullet$ Robustness: To assess robustness under realistic noisy conditions, we compare the mean squared error (MSE) on clean test inputs with that on corrupted inputs containing randomly missing observations, additive perturbations, and occasional demand bursts, and report the performance drop $\Delta\text{MSE}$.
VI-A2 System-Level Benchmarks and Metrics
We then compare FUSION with several benchmark methods enumerated below:
@@ VI-A2 System-Level Benchmarks and Metrics
$\bullet$ Pure online scheduling (PurOnline)[13]: This baseline removes the offline stage of FUSION completely, e.g., ESs and APs do not sign any offline contracts, and UAV locations are not pre-planned. All HUs and MUs directly enter the online market, where their tasks are scheduled only via the PG-BRD-based scheduling over the current ES/AP network environment.
@@
$\bullet$ Two-stage resource provisioning without PG-based online scheduling (FUSION_NoPG): This variant keeps the same offline stage as FUSION, i.e., Off-AIC2 with eACO-VRP is still used to form offline AP–ES contracts and pre-planned routes. However, the online stage abandons PG-BRD: HU/MU tasks are randomly assigned to feasible ESs/APs within coverage subject to resource capacity and deadline constraints.
@@
$\bullet$ Two-stage resource provisioning with random offline contracts (FUSION_Random): This baseline uses the same online stage as FUSION and still relies on PG-BRD for congestion-aware online resource scheduling. However, its offline stage no longer uses Off-AIC2 or eACO-VRP; instead, each AP is randomly associated with a subset of ESs and moves along a random feasible route until its capacity is exhausted.
@@
$\bullet$ Two-stage resource provisioning without ES spatio-temporal features (FUSION_NoST): This baseline preserves the two-stage structure of FUSION, but its offline auction ignores ES-specific time windows and travel-time constraints. Specifically, AP-ES pairs are determined only by simple price/volume trading without using eACO-VRP, while online stage still runs PG-BRD for task scheduling.
@@
TABLE II: Robustness–complexity comparison of different forecasters on synthetic non-stationary series. The “corrupt” case injects 30% random missing values, Gaussian noise with standard deviation $\sigma=0.05$, and random burst shocks.
@@
To assess performance, we use the following metrics:
$\bullet$ Social welfare (SW): SW captures the sum of utilities of all ESs, APs, and users across both stages: in the offline stage, SW accounts for ES profits, AP utilities, and auctioneer surplus given by (1)-(3); in the online stage, it captures the potential value $\Phi(\bm{\pi})$, combining SD-side benefits and SP-side costs.
$\bullet$ Delay incurred by interactions between SPs and SDs (DoI): DoI captures the total decision-making latency (in milliseconds) induced by control-message exchanges between ESs/APs (SPs) and HUs/MUs (SDs) across both offline and online stages. To compute DoI, following [30, 32, 33], we assume that the per-interaction delay lies in the range of $[1,15]$ ms.
$\bullet$ Energy consumption incurred by interactions between SPs and SDs (ECoI): ECoI measures the signaling-related energy consumption during the decision-making process. To compute ECoI, we assume that the transmit power of ESs lies in $[6,20]$ W, consistent with [30, 32]. Specifically, ECoI is obtained by aggregating the energy consumed by all control-message transmissions and the corresponding interaction delays.
$\bullet$ Truthfulness and individual rationality of Off-AIC2: To empirically verify the economic properties of Off-AIC2, we conduct a set of experiments in which ESs and APs are allowed to deviate from their truthful bids/asks. We measure the proportion of ESs and APs whose utilities increase under such deviations. An auction that is truthful and individually rational should ensure that (i) every winning ES attains a non-negative utility and cannot benefit from misreporting its bid, and (ii) every winning AP receives a payment no lower than its declared ask and gains no advantage by inflating or deflating it.
VI-B Performance Evaluation on Prediction
VI-B1 Robustness vs. Cost
We first conduct experiments to evaluate “robustness vs. cost” trade-off in Table II. In the “clean” setting, we feed each model with intact historical sequences, while in the “corrupt” setting, we inject 30% random missing values, additive Gaussian noise with standard deviation $\sigma=0.05$, and random burst shocks of length 1–3. On clean data, LSTM attains the lowest MSE (0.119), whereas Transformer and Pro-LNN yield slightly larger errors (0.467 and 0.459, respectively). This gap is expected, as LSTM uses approximately 52.8k parameters, compared to only 11.1k parameters of Pro-LNN, implying significantly higher computation costs. Once corrupted inputs are introduced, however, LSTM error surges to 0.983 (a relative increase of +726.3%), while the errors of Transformer and Pro-LNN only grow to 0.781 and 0.762 (relative increases of +67.3% and +66.1%), respectively. These results indicate that our proposed Pro-LNN offers a favorable trade-off between robustness and cost compared to an over-parameterized LSTM on noisy sequences, which matches the characteristics of ES demand trajectories in Off-AIC2.
VI-B2 Real-World Dataset Test
[…]
TABLE III: Prediction accuracy and model footprint on the UCI dataset (feeders MT_005 and MT_145, input length $T_{\mathrm{his}}=48$, prediction horizon $H=10$).